# Formalizing Complex Mathematical Statements with LLMs: A Study on Mathematical Definitions

### v2

@@ Abstract
Thanks to their linguistic capabilities, LLMs offer an opportunity to bridge the gap between informal mathematics and formal languages through autoformalization. However, it is still unclear how well LLMs generalize to sophisticated and naturally occurring mathematical statements. To address this gap, we investigate the task of autoformalizing real-world mathematical definitions – a critical component of mathematical discourse. Specifically, we introduce two novel resources for autoformalisation, collecting definitions from Wikipedia (Def_Wiki) and arXiv papers (Def_ArXiv). We then systematically evaluate a range of LLMs, analyzing their ability to formalize definitions into Isabelle/HOL. Furthermore, we investigate strategies to enhance LLMs’ performance including refinement through external feedback from Proof Assistants, and formal definition grounding, where we guide LLMs through relevant contextual elements from formal mathematical libraries. Our findings reveal that definitions present a greater challenge compared to existing benchmarks, such as miniF2F. In particular, we found that LLMs still struggle with self-correction, and aligning with relevant mathematical libraries. At the same time, structured refinement methods and definition grounding strategies yield notable improvements of up to 16% on self-correction capabilities and 43% on the reduction of undefined errors, highlighting promising directions for enhancing LLM-based autoformalization in real-world scenarios. [^1] Code and datasets are available at https://github.com/lanzhang128/definition_autoformalization
Formalizing Complex Mathematical Statements with LLMs:
A Study on Mathematical Definitions
Lan Zhang1, Marco Valentino2, André Freitas1,2,3 1Department of Computer Science, University of Manchester, United Kingdom 2Idiap Research Institute, Switzerland 3National Biomarker Centre, CRUK Manchester Institute, United Kingdom lan.zhang-6@postgrad.manchester.ac.uk {marco.valentino, andre.freitas}@idiap.ch
@@ 1 Introduction
Consequently, one promising application of LLMs is autoformalization, the task of translating informal statements into formal languages (Wu et al., 2022). Given their advanced linguistic capabilities, LLMs offer an opportunity to bridge the gap between informal mathematics, natural language, and machine-verifiable logic, potentially streamlining and scaling the process of formal mathematical reasoning (Jiang et al., 2023; Tarrach et al., 2024).
The task of autoformalization has garnered increasing attention in recent years, leading to the development of benchmarks and evaluation methodologies (Azerbayev et al., 2023; Zhang et al., 2024; Li et al., 2024). Despite this progress, however, existing benchmarks for autoformalization often focus on relatively simple mathematical problems, limiting our understanding of how well LLMs generalize to more sophisticated and naturally occurring mathematical statements.
To address this gap, this paper investigates the task of autoformalizing mathematical definitions – a critical component of mathematical discourse (Moschkovich, 2003). Definitions serve as foundational building blocks in mathematical reasoning, yet they are often intricate, context-dependent, and difficult to formalize. Evaluating LLMs on this subset of mathematical statements, therefore, allows for assessing their ability to formally represent fine-grained mathematical concepts, highlighting persisting challenges and limitations for real-world applications.
@@
Furthermore, to address observed limitations, we investigate two key strategies to enhance performance: (1) Refinement via external feedback, investigating the self-correction capabilities of LLMs by incorporating errors found by the supporting Proof Assistant. In particular, we show that while LLMs exhibit limited ability to refine outputs based on binary feedback (error vs. non-error), a more structured categorical refinement implemented via additional instructional constraints can improve performance. (2) Formal definition grounding. Many mathematical definitions require references to formal objects in external mathematical libraries. To guide LLMs in the process of autoformalization, we investigate the impact of introducing additional contextual control mechanisms, which add contextual elements from formal mathematical libraries as auxiliary premises.
Overall, our findings reveal that the proposed benchmarks present a greater challenge compared to existing autoformalization datasets, such as miniF2F (Zheng et al., 2022). In particular, LLMs struggle with self-correction and particularly with incorporating relevant mathematical libraries as preambles. Proposed structured refinement methods and definition grounding strategies both yield notable improvements, highlighting promising directions for enhancing LLM-based autoformalization in real-world scenarios.
@@
    We perform a comprehensive error analysis on Isabelle/HOL, identifying key failures in formalizations generated by a range of LLMs, including GPT-4o (OpenAI, 2024b), Llama3 (Llama Team, 2024) and DeepSeekMath (Shao et al., 2024).
@@
    We explore the role of formal definition grounding, investigating how the inclusion of relevant mathematical libraries impacts the ability of LLMs to connect the formalized statements with contextual mathematical elements and relevant premises. We found that definition grounding is fundamental for complex autoformalization.
@@
4. The expression $2410x^{2}-x-$ 24 can be written as $(Ax-8)(Bx+3),$ where $A$ and $B$ are integers. What is $AB+B$? Show that it is 12.
@@ 2 Autoformalization with LLMs
Property miniF2F-Test Def_Wiki Def_ArXiv
No. Samples 244 56 30
No. Tokens 70.25(47.70) 200.18(112.98) 164.40(71.47)
No. Objects 4.76(1.68) 7.63(2.71) 7.10(2.64)
No. Formulae 2.71(1.74) 2.84(2.05) 3.17(1.97)
@@ 2.1 Limitations of Existing Benchmarks
Naturally occurring mathematical statements typically involve complex and abstract mathematical concepts. However, the statements in existing datasets, such as miniF2F (Zheng et al., 2022), primarily consist of basic arithmetic operations and elementary mathematical objects, such as integers, fractions, and real numbers (as shown in Table 1). Such mathematical objects are relatively simple compared to the complex and abstract concepts found in naturally occurring mathematical statements and scientific papers, which may involve higher-level structures like vectors, matrices, and probability. The operations are also limited to simple arithmetic, such as addition, subtraction, multiplication, division, and exponentiation. Studying autoformalization on such datasets, therefore, does not necessarily reflect the challenges of autoformalization in realistic scenarios.
In addition, the ground-truth formal code in publicly available datasets may have been exposed to LLMs whose training corpora are not disclosed. Fundamental mathematical definitions also have a high likelihood of already being formalized in theorem prover libraries. This risk raises concerns about data leakage when LLMs perform autoformalization and could lead to biased results when analyzing performance. However, few benchmarks focus on complex mathematical definitions. To address this, we propose constructing data samples of mathematical definitions in the machine learning domain, as concepts in this area are sufficiently complex and less likely to have been formalized.
2.2 Mathematical Definitions in Machine Learning Domain
We obtain mathematical definitions in machine learning domain from two sources: Wikipedia (Def_Wiki) and Arxiv Papers (Def_ArXiv). Definitions from these two sources are likely to have already been validated and exhibit sufficient variety. For Def_Wiki, definitions are from pages under the Machine Learning category [^2] https://en.wikipedia.org/wiki/Category:Machine_learning and its sub-categories. We manually browsed each page, identified well-defined definitions (i.e., formal descriptions with mathematical symbols), and converted the chosen definitions into LaTeX format. In total, we obtained 56 qualified natural language definitions in LaTeX and divided them into development and test sets, containing 10 and 46 samples, respectively. For Def_ArXiv, we used the advanced search tool on ArXiv’s website, filtering for papers in the subcategories cs.LG and stat.ML, with comments including "ICML." We restricted the search to papers published in 2019, 2020, and 2021 and manually reviewed the first 25 papers from each year. We shortlisted papers that were accepted to the ICML conference and contained formally described definitions with mathematical symbols to ensure reliability. We then filtered out definitions that were less straightforward or formal in their expressions, extracted the LaTeX versions, and ultimately obtained 30 definitions from 7 papers.
@@ 2.2 Mathematical Definitions in Machine Learning Domain
Statements in definition datasets are more abstract and complex, as intuitively shown in the randomly chosen examples in Table 1. The data properties are summarized in Table 2. Although small in scale, definition datasets exhibit higher means for the number of tokens, mathematical objects, and formulae per example, indicating that they are more challenging. Additionally, definition datasets have higher standard deviations, suggesting greater diversity among the samples.
@@
The data samples in our benchmarks contain only definitions in LaTeX format. We did not include ground-truth formal codes for the following reasons: 1. Including ground-truth formal codes could increase the risk of the aforementioned data leakage problem. 2. A single mathematical statement can have multiple correct formalizations. An autoformalized code that differs from the ground-truth does not necessarily indicate incorrect formalization. 3. The purpose of ground-truth formal codes is to evaluate autoformalization. However, the syntactic correctness of formalized code can be rigorously and automatically verified using theorem provers (Zhang et al., 2024), and semantic consistency can be evaluated in a reference-free manner (Li et al., 2024). Manual inspection of autoformalized code also does not require ground-truth formal codes.
@@
Table 3: Autoformalization results. Prompt strategies include: (ZS): zero-shot autoformalization; (ZS + Binary): refinement given (zero-shot) formalized code and binary syntactic correctness state. Pass rate (Pass), the place of first error occurrence in the main body of the code (FEO), and percentage of occurrence of each error category are recorded here. Errors in each error category are: (TRO): Time Run-Out for checking; (IVI): Fake Non-Existent Theory, Invalid structural format; (SYN): Inner syntax error, Outer syntax error, Inner lexical error, Malformed command syntax, Bad name, Bad number of arguments for type constructor, Extra free type variable(s); (UDF): Undefined type names, Undeclared class, Undefined locale, No type arity list, Extra variables on rhs; (TUF) Type unification failed.
@@ 3 Empirical Evaluation
Isabelle/HOL is chosen as the representative formal language due to its wide adoption and support for formal mathematical reasoning. We evaluate three LLMs with different features: DeepSeekMath-7B (Shao et al., 2024), Llama3-8B (Llama Team, 2024) and GPT-4o (OpenAI, 2024b). DeepSeekMath-7B is an open-sourced LLM trained specifically for mathematics using mathematical contents from Common Crawl. As a smaller model, it has demonstrated comparable mathematical reasoning performance as in GPT-4 (OpenAI, 2024a), and strong few-shot autoformalization performance on miniF2F with Isabelle. This superiority makes it a good representative of smaller but specialized LLMs. LLama3-8B is a smaller open-sourced foundation LLM with no specific emphasis on math. GPT-4o is widely acknowledged as one of the state-of-the-art LLMs. For reproducibility, greedy decoding is used for generation in all settings.
@@
The success rate of passing the check by the Isabelle Proof Assistant across the tested dataset is used as the first metric. We assume that a formalized code instance with the first error occurring later in the code reflects a more complete autoformalization. Thus, we evaluate such by calculating the proportion of correct lines (up to the first error) within the main body of the code. For syntactically correct instances, this value is equal to 1. To better monitor the occurrence of errors, we group them into three categories: Syntax Errors (SYN), Undefined Item Errors (UDF), and Type Unification Failed Errors (TUF). For each category, we calculate the percentage of incorrect formalized codes caused by errors in that category.
@@ 3.1 Zero-Shot Prompting & Binary Refinement
In order to understand the challenges in autoformalising mathematical definitions with LLMs, we perform a preliminary analysis on miniF2F (Zheng et al., 2022), Def_Wiki and Def_ArXiv using zero-shot prompting (ZS) and binary refinement. With binary refinement, we aim to assess the capabilities of LLMs for error correction, providing them with the formal code generated via ZS, along with the syntactic correctness evaluated using the proof assistant (i.e., “correct”, “incorrect”). From the results reported in Table 3, we can derive the following observations:
Def_Wiki and Def_ArXiv are more challenging than miniF2F.
@@
In Table 3, we observe that the percentage of Invalid Inputs errors (IVI) can be non-zero. Errors in this category are caused by either non-existent preambles or invalid theory file formats in structure. For Llama3-8B the latter is more common whereas for GPT-4o, we observe that the dominant cause is the generation of non-existent preambles. This behaviour shows that LLMs do not perfectly generalize in recognizing the names of preambles.
@@
As a model designed specifically for mathematics, DeepSeekMath with 7B parameters can achieve a similar success rate as GPT-4o. Although Llama3-8B has a larger model size, its generalization ability on definitions is limited. Additionally, DeepSeekMath-7B exhibits a lower percentage of undefined type names errors (UDF). However, one disadvantage of the specialized model is that its formalizations have a higher percentage of time run-out issues (TRO). This is likely caused by the bias introduced during the fine-tuning phase on theorem proving which can lead the model to generate unsolicited proofs.

### v3

# Autoformalization in the Wild: Assessing LLMs on Real-World Mathematical Definitions

@@ Abstract
Thanks to their linguistic capabilities, LLMs offer an opportunity to bridge the gap between informal mathematics and formal languages through autoformalization. However, it is still unclear how well LLMs generalize to sophisticated and naturally occurring mathematical statements. To address this gap, we investigate the task of autoformalizing real-world mathematical definitions: a critical component of mathematical discourse. Specifically, we introduce two novel resources for autoformalization, collecting definitions from Wikipedia (Def_Wiki) and arXiv papers (Def_ArXiv). We then systematically evaluate a range of LLMs, analyzing their ability to formalize definitions into Isabelle/HOL. Furthermore, we investigate strategies to enhance LLMs’ performance including refinement through external feedback from Proof Assistants, and formal definition grounding, where we augment LLMs’ formalizations through relevant contextual elements from formal mathematical libraries. Our findings reveal that definitions present a greater challenge compared to existing benchmarks, such as miniF2F. In particular, we found that LLMs still struggle with self-correction, and aligning with relevant mathematical libraries. At the same time, structured refinement methods and definition grounding strategies yield notable improvements of up to 16% on self-correction capabilities and 43% on the reduction of undefined errors, highlighting promising directions for enhancing LLM-based autoformalization in real-world scenarios. [^1] Code and datasets are available at https://github.com/lanzhang128/definition_autoformalization
Autoformalization in the Wild: Assessing LLMs on
Real-World Mathematical Definitions
Lan Zhang1, Marco Valentino2, André Freitas1,3,4 1Department of Computer Science, University of Manchester, United Kingdom 2School of Computer Science, University of Sheffield, United Kingdom 3Idiap Research Institute, Switzerland 4National Biomarker Centre, CRUK Manchester Institute, United Kingdom lan.zhang-6@postgrad.manchester.ac.uk m.valentino@sheffield.ac.ukandre.freitas@idiap.ch
@@ 1 Introduction
Consequently, one promising application of LLMs is autoformalization, the task of translating informal statements into formal languages (Wu et al., 2022). Given their advanced linguistic and inferential capabilities, LLMs offer an opportunity to bridge the gap between informal mathematics, natural language, and machine-verifiable logic, potentially streamlining and scaling the process of formal mathematical reasoning (Jiang et al., 2023; Tarrach et al., 2024).
The task of autoformalization has collected increasing attention in recent years, leading to the development of benchmarks and evaluation methodologies (Azerbayev et al., 2023; Zhang et al., 2024; Li et al., 2024). Despite this progress, however, existing benchmarks for autoformalization often focus on relatively simple mathematical problems, limiting our understanding of how well LLMs generalize to more sophisticated and naturally occurring mathematical statements.
To address this gap, this paper investigates the task of autoformalizing mathematical definitions, a critical component of mathematical discourse (Moschkovich, 2003). Definitions serve as foundational building blocks in mathematical reasoning, yet they are often intricate, context-dependent, and thus difficult to formalize. Evaluating LLMs on this subset of mathematical statements, therefore, allows for assessing their ability to formally represent fine-grained mathematical concepts, highlighting persisting challenges and limitations for real-world applications.
@@
Furthermore, to address observed limitations, we investigate two key strategies to enhance performance: (1) Refinement via external feedback, investigating the self-correction capabilities of LLMs by incorporating errors found by the supporting Proof Assistant. In particular, we show that while LLMs exhibit limited ability to refine outputs based on binary feedback (error vs. non-error), a more structured categorical refinement implemented via additional instructional constraints can improve performance. (2) Formal definition grounding. Many mathematical definitions require references to formal objects in external mathematical libraries. To augment autoformalization from LLMs, we investigate the impact of introducing additional contextual control mechanisms, which add contextual elements from formal mathematical libraries as auxiliary premises.
Overall, our findings reveal that the proposed benchmarks present a greater challenge compared to existing autoformalization datasets, such as miniF2F (Zheng et al., 2022). In particular, LLMs struggle with self-correction and particularly with incorporating relevant mathematical libraries as preambles. Proposed structured refinement methods and definition grounding strategies both deliver notable improvements, highlighting promising directions for enhancing LLM-based autoformalization in real-world scenarios.
@@
    We perform a comprehensive error analysis on Isabelle/HOL, identifying key failures in formalizations generated by LLMs spanning across different families, including GPT-4o (OpenAI, 2024b), Llama3 (Llama Team, 2024) and DeepSeekMath (Shao et al., 2024).
@@
    We explore the role of formal definition grounding, investigating how the inclusion of relevant mathematical libraries impacts the ability of LLMs to connect the formalized statements with contextual mathematical elements and relevant premises.
@@
4. The expression $10x^{2}-x-24$ can be written as (Ax8)(Bx+3),(Ax-8)(Bx+3), where $A$ and $B$ are integers. What is $AB+B$? Show that it is 12.
@@ 2.1 Limitations of Existing Benchmarks
Naturally occurring mathematical statements typically involve complex and abstract mathematical concepts. However, the statements in existing datasets, such as miniF2F (Zheng et al., 2022), primarily consist of basic arithmetic operations and elementary mathematical objects, such as integers, fractions, and real numbers (as shown in Table 1). Such mathematical objects are relatively simple compared to the complex and abstract concepts found in naturally occurring mathematical statements and scientific papers, which may involve higher-level structures like vectors, matrices, and probability. The operations are also limited to simple arithmetic, such as addition, subtraction, multiplication, division, and exponentiation. Studying autoformalization on such datasets, therefore, does not necessarily reflect the challenges of autoformalization in realistic scenarios. However, few benchmarks focus on how to construct complex mathematical statements. Our work aims to address this gap.
2.2 Real-World Mathematical Definitions
Since extracting high-quality definition statements from general mathematical corpora requires careful curation, we propose a systematic data creation process that balances complexity and diversity. We begin by classifying real-world definitions into two categories: (i) common definitions, which are presented in a global context where the necessary preliminaries are implicit and relatively general, and (ii) specialized definitions, which are typically situated in a local context and rely on explicit and specific preliminaries. Both types present distinct challenges for autoformalization. In the former, the model must infer implicit preliminaries and connect them with existing formal constructs. In the latter, the model must not only formalize the definition but also the associated preliminaries. We ground these two types of definitions in two sources: Wikipedia for common definitions (Def_Wiki) and Arxiv Papers for specialized definitions (Def_ArXiv), as definitions from these two sources are human-written, naturally occurring and likely to have already been validated.
To construct a representative instance for Def_Wiki and Def_Arxiv, we focus on definitions from the field of machine learning, as this domain offers a sufficient number of novel, diverse, and relatively unformalized definitions. Moreover, the task of formalizing such definitions holds practical value for the AI community. The detailed operational steps are provided in Appendix A. A quality check was performed to ensure that the selected definitions exhibit high diversity and present complementary challenges. This process yielded 56 definitions for Def_Wiki and 30 definitions for Def_Arxiv. This number is expected, as novel definitions in the real world are relatively scarce compared to example questions in mathematics (such as those in miniF2F), which are often closer in nature to synthetically generated content. Although the resulting datasets are relatively small in scale, they are sufficient to expose the core challenges of autoformalization in real-world scenarios. The size of the datasets is not the primary factor determining the significance, transferability, or robustness of the findings. Exploring additional scientific domains and expanding the dataset further are promising directions for future work.
@@ 2.2 Real-World Mathematical Definitions
We compare the quality of miniF2F with the target definition datasets. MiniF2F is significantly less abstract, complex and diverse, as intuitively shown in the randomly chosen examples in Table 1. The data properties are summarized in Table 5 in Appendix. Definition datasets exhibit higher means for the number of tokens, mathematical objects, and formulae per example, indicating that they are significantly more complex. Additionally, definition datasets have higher standard deviations, suggesting greater diversity among samples.
@@
The proposed benchmarks contain only definitions in LaTeX format. We did not include ground-truth formal code for the following reasons: 1. Including such code could increase the risk of data leakage, as ground-truth formalizations from publicly available datasets may have been seen by LLMs, whose training data is not fully disclosed, making it difficult to determine whether future improvements genuinely address the challenges posed by our benchmark or simply reflect prior exposure. 2. A single mathematical statement can have multiple correct formalizations. An autoformalized output that differs from the reference does not necessarily indicate incorrect formalization. 3. The main purpose of ground-truth formal code is to evaluate autoformalization. However, the syntactic correctness of formalized code can be rigorously and automatically verified using theorem provers (Zhang et al., 2024), while semantic consistency can potentially be assessed in a reference-free manner via LLM-as-Judges (Zhang et al., 2025). Moreover, manual inspection of autoformalized code does not require ground-truth formalizations.
@@
Table 2: Autoformalization results. Prompt strategies include: (ZS): zero-shot prompting; (ZS + Binary): refinement given (ZS) formalized code and binary syntactic correctness state. Pass rate (Pass), the place of first error occurrence in the main body of the code (FEO), and percentage of occurrence of each error category are recorded here. Errors in each error category are: (TRO): Time Run-Out for checking; (IVI): Fake Non-Existent Theory, Invalid structural format; (SYN): Inner syntax error, Outer syntax error, Inner lexical error, Malformed command syntax, Bad name, Bad number of arguments for type constructor, Extra free type variable(s); (UDF): Undefined type names, Undeclared class, Undefined locale, No type arity list, Extra variables on rhs; (TUF) Type unification failed.
@@ 3 Empirical Evaluation
Isabelle/HOL was chosen as the representative formal language due to its widespread adoption within the formal mathematics community and its ability to provide specific information about the types of errors encountered. Moreover, Isabelle employs a declarative proof language that is closer to structured natural language, making it well-suited for exploring autoformalization of complex natural language statements. We prioritize an in-depth analysis of a single formal language over a broad comparison across multiple languages. We believe this approach lays the foundation for future work aimed at exploring the behavior of LLMs in alternative formal systems. We evaluate three LLMs with different features: DeepSeekMath-7B (Shao et al., 2024), Llama3-8B (Llama Team, 2024) and GPT-4o (OpenAI, 2024b). DeepSeekMath-7B is an open-sourced LLM trained specifically for mathematics. As a smaller model, it has demonstrated comparable mathematical reasoning performance as in GPT-4 (OpenAI, 2024a), and strong few-shot autoformalization performance on miniF2F with Isabelle. This superiority makes it a good representative of smaller but specialized LLMs. LLama3-8B is a smaller open-sourced foundation LLM with no specific emphasis on math. GPT-4o is widely acknowledged as one of the state-of-the-art LLMs. For reproducibility, greedy decoding is used for generation in all settings.
@@
The success rate of passing the check by the Isabelle Proof Assistant across the tested dataset is used as the first metric. We assume that a formalized code instance with the first error occurring later in the code reflects, as a proxy, the level of autoformalization. Thus, we evaluate such by calculating the proportion of correct lines (up to the first error) within the main body of the code. For syntactically correct instances, this value is equal to 1. To better monitor the occurrence of errors, we group them into three categories: Syntax Errors (SYN), Undefined Item Errors (UDF), and Type Unification Failed Errors (TUF). For each category, we calculate the percentage of incorrect formalized codes caused by errors in that category.
@@ 3.1 Zero-Shot Prompting & Binary Refinement
In order to understand the challenges in autoformalizing mathematical definitions with LLMs, we perform a preliminary analysis on miniF2F (Zheng et al., 2022), Def_Wiki and Def_ArXiv using zero-shot prompting (ZS) and binary refinement. With binary refinement, we aim to assess the capabilities of LLMs for error correction by providing them with the formal code generated via ZS, along with the syntactic correctness evaluated using the proof assistant (i.e., “correct”, “incorrect”). From the results reported in Table 2, we can derive the following observations:
Def_Wiki and Def_ArXiv are significantly more challenging than miniF2F.
@@
In Table 2, the percentage of Invalid Inputs errors (IVI) can be non-zero. Errors in this category are caused by either non-existent preambles or invalid file formats in structure. For Llama3-8B the latter is more common whereas for GPT-4o, we observe that the dominant cause is the generation of non-existent preambles, showing that GPT-4o do not perfectly generalize in recognizing the names of preambles.
Category Reasons
UDF 1. Items not defined. Formalization requires every mentioned item to be clearly defined in the local context or preambles. During autoformalization, GPT-4o could refer to items that are not defined in both sources.
@@
As a model designed specifically for mathematics, DeepSeekMath-7B can achieve a similar success rate as GPT-4o. Although Llama3-8B has a larger model size, its generalization ability on definitions is limited. Additionally, DeepSeekMath-7B exhibits a lower percentage of undefined type names errors (UDF). However, one disadvantage of the specialized model is that its formalizations have a higher percentage of time run-out issues (TRO). This is likely caused by the bias introduced during the fine-tuning phase on theorem proving which can lead the model to generate unsolicited proofs.
@@ 3.1.1 Error Analysis & Interventions
We observe that syntactic errors (SYN) exhibit the most variety, suggesting that GPT-4o may struggle to follow syntactic rules in Isabelle/HOL if not explicitly instructed. Type unification errors (TUF) suggest that GPT-4o may struggle with the exact usage of defined Isabelle items. To improve these issues, we investigate a Categorical Refinement (CR) method. CR involves designing specific additive instructions that constraint the behaviors leading to errors identified in the qualitative analysis.
Similarly, for syntactic errors (SYN), causes 1, 2, and 3 in Table 3 can be addressed with rule-based algorithms that refine formal codes at the symbolic level (Symbolic Refinement, SR). Undefined errors (UDF), on the other hand, indicate that although GPT-4o can refer to external formal mathematical items, it remains unaware of the location or existence of relevant libraries. To alleviate UDF errors, we propose the process of Formal Definition Grounding (FDG): linking mathematical objects mentioned in natural language statements to their formal definitions in formal libraries, and incorporating this information as contextual elements for formalizations.
3.2 Categorical Refinement